# AnimateScene: Camera-controllable Animation in Any Scene

### v1

@@ 4.1 Experimental setup
Datasets. We utilize 83 single-view, copyright-free online photos and 83 generated 4D Human animation samples for evaluation. The image set encompasses both indoor and outdoor environments, spanning real-world and simulated scenes. The 4D Human data features a diverse range of dance and sports movements generated by LHM (Qiu et al. 2025). Each 4D Human instance is paired with a carefully selected scene to best complement its motion characteristics. We designed six types of trajectories to evaluate the quality of scene generation with controllable cameras. More details can be found in the supplementary material.
Evaluation metrics. Drawing inspiration from previous works (Wang et al. 2024b; Wang, Chan, and Loy 2023), we adopt the Vision-Language Model LLaVA (Liu et al. 2023) to evaluate the rendered synthetic scene video frames with different camera trajectories. The LLaVA-IQA metric focuses on five dimensions: noise suppression, sharpness of edges, structural consistency, level of detail, and overall perceptual quality. Details of the LLaVA-IQA metric are given in supplementary.
Baselines. Since there is currently no method for reconstructing camera controllable 3D background scenes for 4D humans, we compare our method with one 3D scene reconstruction work 3DitScene (Zhang et al. 2024) and two generative view synthesis method SEVA (Zhou et al. 2025b) and DimensionX (Sun et al. 2024). 3DitScene is a unified scene editing framework leveraging language-guided disentangled gaussian splatting that enables seamless editing from 2D to 3D, allowing generate 3D scene from single image and synthesize novel viewpoints. SEVA and DimensionX are novel view video generation method. SEVA is a generalist diffusion model that creates novel views of a scene, given any number of input views and target cameras. DimensionX is a framework designed to generate photorealistic 3D and 4D scenes from just a single image with video diffusion.
@@ 4.2 Evaluation results
Method Noise-Free $\uparrow$ Edge $\uparrow$ Sturcture $\uparrow$ Detail $\uparrow$ Quality $\uparrow$
3DitScene (Zhang et al. 2024) 0.696 0.262 0.533 0.645 0.538
SEVA (Zhou et al. 2025b) 0.432 0.045 0.291 0.429 0.312
DimensionX (Sun et al. 2024) 0.759 0.135 0.392 0.817 0.559
@@
Table 1: LLaVA-IQA scores on synthetic video frames. AnimateScene ranks highest on five metrics (noise suppression, edge sharpness, structural consistency, detail level, and overall perceptual quality).
3DitScene (Zhang et al. 2024) SEVA (Zhou et al. 2025b) DimensionX (Sun et al. 2024) Ours
@@
Table 2: The percentage of favorite users for the video quality and geometric consistency of video generated by each method. Human and GPT-4v vote shares for each method’s video quality and geometric consistency, with AnimateScene leading both.
@@
Quantitative results. We provide a project page (HTML) in the supplementary materials, a local HTML file (project_page.html), which exhibits the videos generated by our method and baselines. To view the qualitative results, please open project_page/project_page.html in a web browser. Since there is currently no method for reconstructing camera controllable 3D background scenes for 4D humans, we composite the baseline-generated controllable-view 3D scenes with the 4D human sequences frame by frame. Specifically, we render the human sequences according to the predefined camera trajectory, and then paste each rendered human frame onto the baseline-generated background according to positions provided by the object placement model, which ensures geometric consistency between the human and the scene in the final video. In Table 1, we report the LLaVA scores across different dimensions, which showed that our method is capable of synthesizing high-quality scenes with controllable camera trajectories.
@@
To further evaluate the synthetic result, we conduct a user study to compare the rendered video by our method with the established baselines. We ask 20 users to vote for their preferred method based on video quality and geometric consistency across different viewpoints. Followed 3DitScene (Zhang et al. 2024), we also adopt GPT-4v as an additional criterion since it can evaluate 3D consistency and image quality well aligned with human preference. In Table 2, we report the user study result, which shows our method outperforms other baselines in terms of both video quality and geometric consistency.
Figure 4: Ablation study on object placement and scene reconstruction. Each row shows the scene (col. 1), human (col. 2), baseline without key components (col. 3), and our full AnimateScene (col. 4). In rows 1–2, 2D placement yields a sinking foot or floating figure, whereas ours produces stable, natural poses. In rows 3–4, skipping view‐fusion inpainting and post‐processing leaves black holes by the sofa and road; our full method fills them for a coherent 3D scene.
Qualitative results. In Figure 3, we show qualitative results of our methods and other baselines. Since different reconstruction can render different view under the same camera position, and the positional information is encoded as a condition in video generation models, the corresponding frames produced by each method still differ slightly even we use the same camera trajectory. It can be seen that 3DitScene (Zhang et al. 2024) produces noticeable artifacts when the camera rotation leads to a significant deviation from the original view. In contrast, video generation methods like SEVA (Zhou et al. 2025b) and DimensionX (Sun et al. 2024) lack geometric consistency. Moreover, as these methods are not end-to-end frameworks, they cannot model the geometric interaction between humans and the scene, resulting in unrealistic compositions. In comparison, our method outperforms others in both geometric consistency and image quality. This demonstrates that our approach can provide camera controllable backgrounds for 4D human sequences within a fully automated synthesis pipeline.
4.3 Ablation study
We ablate key modules in Fig. 4. Replacing our 3D placement network with a naïve 2D depth-sampling baseline (rows 1–2) makes the avatar sink or float, whereas our method yields stable, scene-anchored poses. Substituting our joint post-reconstruction with VistaDream’s pipeline (rows 3–4) leaves black cavities beside the sofa and on the road, but our full model inpaints these gaps for a seamless scene. The qualitative differences are striking even under casual inspection and persist across a wide variety of indoor and outdoor scenes. A quantitative summary in the supplementary material further corroborates these visual findings, confirming that both 3D placement and view-fusion inpainting are critical for high-quality dynamic scene generation.
5 Conclusion and future work
@@
AnimateScene converts a single scene image, human image, motion clip, and camera path into a controllable 4D video. It first matches the actor’s appearance to the scene lighting, then embeds a depth-guided 3D avatar, and finally inpaints view-dependent gaps for seamless composition. Extensive quantitative tests and a user study show AnimateScene surpasses state-of-the-art baselines in noise suppression, edge sharpness, structural consistency, detail, perceptual quality, and geometric coherence. Future work will extend to large-scale outdoor scenes with dynamic lighting and materials, support multi-actor interactions, enable real-time reconstruction and rendering, and incorporate language- and semantics-driven controls for high-level edits and narrative direction.

### v2

@@ 3.1 Experimental setup
We evaluate on 83 copyright-free single-view photos (indoor/outdoor, real and simulated) paired with 83 4D-human animation clips generated by LHM [12]; each human instance is matched to a complementary scene, and six camera trajectories are designed to probe camera-controllable scene generation. Following prior work [15], we assess rendered frames under different camera paths with the VLM LLaVA [8], reporting LLaVA-IQA on noise suppression, edge sharpness, structural consistency, detail, and overall perceptual quality. As no existing method reconstructs camera-controllable 3D backgrounds for 4D humans, we compare against 3DitScene [18]—a language-guided disentangled Gaussian-splatting framework enabling single-image 3D scene generation and novel-view synthesis—and two generative view-synthesis methods, SEVA [20] (a generalist diffusion model for novel views given arbitrary inputs and target cameras) and DimensionX [14] (photorealistic 3D/4D scene generation from a single image via video diffusion).
@@ 3.2 Evaluation results
Method Noise-Free $\uparrow$ Edge $\uparrow$ Structure $\uparrow$ Detail $\uparrow$ Quality $\uparrow$
3DitScene [18] 0.696 0.262 0.533 0.645 0.538
SEVA [20] 0.432 0.045 0.291 0.429 0.312
DimensionX [14] 0.759 0.135 0.392 0.817 0.559
@@
Table 1: LLaVA-IQA scores on synthetic video frames. AnimateScene ranks highest on five metrics.
3DitScene [18] SEVA [20] DimensionX [14] Ours
@@
Table 2: Human and GPT-4V preference shares on video quality and geometric consistency across methods, with AnimateScene achieving the highest votes.
@@
Since there is currently no method for reconstructing camera controllable 3D background scenes for 4D humans, we composite the baseline-generated controllable-view 3D scenes with the 4D human sequences frame by frame. Specifically, we render the human sequences according to the predefined camera trajectory, and then paste each rendered human frame onto the baseline-generated background according to positions provided by the object placement model, which ensures geometric consistency between the human and the scene in the final video. In Table 1, we report the LLaVA scores across different dimensions, which showed that our method is capable of synthesizing high-quality scenes with controllable camera trajectories.
@@
To further evaluate the synthetic result, we conduct a user study to compare the rendered video by our method with the established baselines. We ask 20 users to vote for their preferred method based on video quality and geometric consistency across different viewpoints. Followed 3DitScene [18], we also adopt GPT-4v as an additional criterion since it can evaluate 3D consistency and image quality well aligned with human preference. In Table 2, we report the user study result, which shows our method outperforms other baselines in terms of both video quality and geometric consistency.
4 Conclusion
@@
AnimateScene converts a single scene image, human image, motion clip, and camera path into a controllable 4D video. It first matches the actor’s appearance to the scene lighting, then embeds a depth-guided 3D avatar, and finally inpaints view-dependent gaps for seamless composition. Extensive quantitative tests and a user study show AnimateScene surpasses state-of-the-art baselines in noise suppression, edge sharpness, structural consistency, detail, perceptual quality, and geometric coherence.
5 Acknowledgement
The work was supported by the National Natural Science Foundation of China (Grant No. 62471287).